OpenStudy (anonymous):

The price of a baseball ticket at Yankee Stadium increased by 2 1∕4 over the last three years. If the original price of a ticket was $60, what is the price of the ticket today?

OpenStudy (anonymous):

Let P be the original cost of the ticket. Let CP be the current cost of a ticket (3 years later) CP = P + 2.25*P CP = 3.25*P CP = 3.25 * 60 CP = $195.00

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I got 135 originally because i didnt add the 1.00